Total Cost Method of Product Pricing

Smart Stream Inc. uses the total cost method of applying the cost-plus approach to product pricing. The costs of producing and selling 6,000 units of cell phones are as follows:

Variable costs: Fixed costs:
    Direct materials $ 89 per unit     Factory overhead $270,800
    Direct labor 41     Selling and administrative expenses 95,200
    Factory overhead 27
    Selling and administrative expenses 21
         Total variable cost per unit $178 per unit

Smart Stream desires a profit equal to a 14% return on invested assets of $798,940.

a. Determine the total cost and the total cost amount per unit for the production and sale of 6,000 units of cellular phones. Round the cost per unit to two decimal places.

Total cost $
Total cost amount per unit $

b. Determine the total cost markup percentage (rounded to two decimal places) for cellular phones.
%

c. Determine the selling price of cellular phones. Round to the nearest cent.
$ per cellular phone
Answer #1
a
Total variable cost 1068000 =178*6000
Total Fixed cost 366000 =270800+95200
Total cost 1434000
Total cost amount per unit 239 =1434000/6000
b
Required markup 111851.6 =798940*14%
Divide by Total cost 1434000
Total cost markup percentage 7.80%
c
Total cost amount per unit 239.00
Add: Markup per unit 18.64 =239*7.8%
Selling price 257.64 per cellular phone